# CLIP Meets Diffusion: A Synergistic Approach to Anomaly Detection

### v1

@@ Abstract
Anomaly detection is a complex problem due to the ambiguity in defining anomalies, the diversity of anomaly types (e.g., local and global defect), and the scarcity of training data. As such, it necessitates a comprehensive model capable of capturing both low-level and high-level features, even with limited data. To address this, we propose CLIPFUSION, a method that leverages both discriminative and generative foundation models. Specifically, the CLIP-based discriminative model excels at capturing global features, while the diffusion-based generative model effectively captures local details, creating a synergistic and complementary approach. Notably, we introduce a methodology for utilizing cross-attention maps and feature maps extracted from diffusion models specifically for anomaly detection. Experimental results on benchmark datasets (MVTec-AD, VisA) demonstrate that CLIPFUSION consistently outperforms baseline methods, achieving outstanding performance in both anomaly segmentation and classification. We believe that our method underscores the effectiveness of multi-modal and multi-model fusion in tackling the multifaceted challenges of anomaly detection, providing a scalable solution for real-world applications.
@@ 1 Introduction
Anomaly detection is an important problem that is widely used in various fields including industry [2, 60] and medicine [47, 35]. In this paper, we address anomaly classification and segmentation.111It is also referred to as anomaly detection and localization, respectively, in the literature. We use the terms anomaly classification and segmentation, and refer to anomaly detection as the overarching problem encompassing both. In anomaly classification (or segmentation), each image (or pixel) is classified as normal or anomalous.
@@
In practice, the number of abnormal images is small, while the potential types of abnormalities are highly diverse, making it challenging to obtain representative examples. Moreover, developing methods that can quickly adapt to new environments (e.g., new processes or products), even with a limited number of normal images, has become a critical research topic [44, 48, 60, 23]. To address this, we focus on the problem within the zero-shot or few-normal-shot222In this paper, ”few-shot” specifically refers to ”few-normal-shot.” setup.
@@
The recent advent of large-scale pretrained foundation models, e.g., GPT [4] and DALL-E [38, 39], has opened new horizons for addressing such challenges. The generalizability and scalability of foundation models make them particularly effective for tackling zero-shot and few-shot tasks such as anomaly detection. In this paper, we present a novel methodology that combines the complementary strengths of discriminative and generative foundation models.
@@
Anomaly detection is a multifaceted problem, and various approaches have been developed based on different perspectives. First, anomaly detection has the aspect of one-class classification. This has led the community to methods that learn normality from normal images and judge what deviates from it as anomalous. Here, many methods extract normal features from normal images using a vision foundation model [10, 11, 43]. Second, anomaly detection has the aspect of two-class classification. The community has employed vision-language foundation models, such as CLIP [37], to compare a query image against normal and abnormal text prompts [25, 29]. Third, anomaly detection also has the aspect of semantic segmentation. Many anomalous features are fine-grained ones that appear in textures or edges. WinCLIP (window-based CLIP) attempted to solve this by sliding windows at multiple scales. However, it is known that CLIP has limitations in learning fine-grained features [51, 26], and such an approach might be suboptimal. Also, it requires hundreds of forward passes due to the exhaustive sliding of multi-scale windows.
@@
We claim that the multifaceted nature of anomaly detection makes it challenging to address using only discriminative models. So, we aim to create synergy by fusing a discriminative vision-language model (VLM) with a generative text-to-image (T2I) model. These are two fundamentally distinct approaches to aligning images and text. Discriminative models, such as CLIP, directly align embeddings in the same space, while generative models, like diffusion models, align them indirectly by jointly performing a common task. Diffusion models learn the relationships between images and texts in a more flexible way, enabling them to capture semantic relationships with greater detail. Therefore, it is essential to leverage both high-level information extracted from the CLIP-based model and low-level information extracted from the diffusion-based model for effective anomaly detection.
@@
We propose CLIPFUSION (= CLIP + Diffusion), a novel method that combines the strengths of multi-modal (language and vision) and multi-model (CLIP and diffusion) approaches. In the zero-shot scenario, CLIPFUSION effectively identifies anomalies by integrating image-text alignment from a CLIP-based model and cross-attention maps [7, 55, 33] from a diffusion-based model. In the few-shot scenario, it further enhances its performance by additionally leveraging feature maps from the CLIP image encoder and the diffusion denoiser, enabling a more comprehensive comparison between the query image and reference images. Figure 1 illustrates qualitative results, showcasing the model’s ability to capture fine-grained anomalies. CLIPFUSION exemplifies a holistic framework that synergizes complementary foundation models, and has the potential to extend beyond anomaly detection and advance other recognition tasks.
@@
We introduce a novel use of diffusion models as feature extractors for anomaly detection. To our knowledge, this is the first approach to directly construct anomaly maps using cross-attention and feature maps from the diffusion denoiser, making the process fast despite using a diffusion model.
We introduce a novel approach to anomaly detection that jointly leverages CLIP and diffusion models, going beyond previous one-sided approaches. We demonstrate the complementary effect of combining features from discriminative and generative models. Since diffusion models compensate for CLIP’s limitations in capturing fine-grained local details, our method eliminates the need for complex and time-consuming techniques such as sliding windows.
Our method consistently outperforms existing approaches in anomaly detection benchmarks. Unlike most prior works, our method is versatile: it can handle both segmentation and classification tasks under zero-shot and few-shot settings.
Our method is training-free, which makes it robust and easy to apply. In contrast, training-based methods are sensitive to the choice of when to stop training, which is challenging in few-shot setups due to the absence of a validation set.
@@ 2 Related Work
In works that utilize vision(-language) models in anomaly detection, features are extracted from the intermediate layers of a model pretrained on a discriminative task (e.g., classification or contrastive learning). Anomaly detection is performed by comparing the features of the query image with the features of reference images. Some works utilize ResNet-based models pretrained on ImageNet in the few-shot regime. In SPADE [10], a -nearest neighbors algorithm is employed to identify the reference images closest to the query image, which are then used as the basis for anomaly detection. In PaDiM [11], features extracted from reference images are modeled as Gaussian distributions, and the features of the query image are compared against these distributions. In PatchCore [43], a memory bank is constructed using the features of reference images, and the features of a query image are directly compared against this memory bank. WinCLIP [25] applies local windows and combines the results through harmonic aggregation. PromptAD [29] learns prompt tokens through contrastive training.
A separate line of work assumes access to additional data and leverages it. Some methods require an auxiliary dataset containing both normal and anomaly images, e.g., training on the MVTec-AD dataset and inference on the VisA dataset [22, 59, 52, 32, 5, 58, 16]. AnomalyGPT [17] relies on an auxiliary dataset with a large number of normal samples. In practice, acquiring such data with matching characteristics is challenging, and these methods often struggle when deployed on datasets with distribution shifts. Some methods also rely on ground-truth segmentation masks, which are difficult to obtain in real-world scenarios [22, 5, 58]. MuSc [30] leverages the test set by exhaustively comparing the features of all test images, which is only feasible when the entire test set is available and contains a low proportion of anomalies.
@@
Recently, there has been growing interest in applying diffusion models to discriminative tasks [15]. In DDPM-Seg [1], feature maps are constructed using activations extracted from intermediate layers of the denoiser and are utilized for semantic segmentation. VPD [56] further incorporates cross-attention maps for semantic segmentation and depth estimation. These methods require fine-tuning for downstream tasks. On the other hand, DiffusionClassifier [27] performs zero-shot classification by determining which text conditioning best predicts the noise. P2PCAC [20] and USCSD [19] reveal that cross-attention maps contain rich information about spatial layouts through image editing and semantic correspondence tasks.
@@ 3 Preliminaries
CLIP [37] consists of an image encoder and a text encoder . These encoders map images and texts into the same embedding space. The model is trained to match correct image-text pairs with high similarity while assigning low similarity to incorrect pairs. This process is optimized using the following contrastive loss [8]:
@@
Diffusion Models
Diffusion models [21] are generative models that produce data by reversing a gradual noising process. In the forward process, Gaussian noise is incrementally added to clean data over multiple steps, transforming it into pure noise. The reverse process is learned via a neural network that denoises the data step-by-step to recover the original signal. When a text prompt is provided, it acts as a guide so that the image being generated matches the meaning or description in the text. They can also be utilized for inpainting a specific region of an image. At each step, the untouched region is preserved by replacing it with the corresponding part of the original image, ensuring that only the specified region is generated and seamlessly blended with the rest of the image [40]. See Appendix A for more details.
@@ 4 Method
The final model, CLIPFUSION, is a fusion of our simple CLIP-based model and our complementary diffusion-based model. For convenience, we call the CLIP-based model PatchCLIP (patch-based CLIP) and the diffusion-based model MapDiff (map-based diffusion). We first describe the characteristics of each model and what resources can be extracted from them for later use.
@@
CLIP processes an image by dividing it into multiple patches and outputs both patch embeddings and a class token embedding. While the class token embedding is trained to be aligned with text embeddings, we observe that patch embeddings are also beneficial, so we leverage both. Given that fine-grained details are handled by the diffusion-based model, PatchCLIP provides a simple yet effective spatial prior, avoiding the cumbersome multi-scale window sliding in WinCLIP.
Additionally, we utilize features extracted from CLIP’s image encoder. To achieve this, we make use of the VV-attention mechanism proposed in CLIP surgery [31, 29]. This mechanism was introduced to complement CLIP’s QK-attention mechanism, which is specialized in global feature extraction. It facilitates local feature extraction by enabling direct interactions between values (hence the name VV-attention) that carry information about the patches themselves.
@@
We extract and utilize cross-attention maps and feature maps from the diffusion denoiser. To achieve this, we use an inpainting pipeline that takes a clean image as input, rather than a generation pipeline that starts with random noise. In the inpainting pipeline, a mask is provided to specify the region where noise will be added. Here, we use the trick of applying an empty mask. This prevents noise from being added, and the denoiser perceives the original image as noisy, allowing us to control the timestep to extract features at the desired level.
@@ 4.1.1 Zero-shot
In the zero-shot setup, given a query image, we leverage the relationships between images and texts learned by multi-modal foundation models to create anomaly score maps.
@@
The query image is input to the CLIP image encoder (e.g., ViT), which divides it into patches, processes them, and outputs patch embeddings. On the other hand, normal and abnormal text prompts are input to the CLIP text encoder and normal and abnormal text embeddings are output. Then, the degree (Equation 1) to which each patch embedding is aligned to the abnormal text embedding is calculated for all patches and organized into a map. This is upsampled to the original image size through interpolation to obtain .
@@
where [object] refers to the entity (e.g., “bottle") in the image, and [state] describes a possible abnormal condition (e.g., “crack") of it. The [state] token interacts with the image as it passes through the denoiser (e.g., U-Net), generating a cross-attention map. A higher value at a specific location in the map indicates that the location responds more strongly to the [state] token (i.e., the abnormal condition). Therefore, we can use this cross-attention map as an anomaly score map (Figure 3). Additionally, we calculate the average of cross-attention maps derived using various possible abnormal conditions (e.g., “crack", “hole", “residue", “damage") to obtain .
@@
where represents the set of text prompts that describe abnormal conditions. Assuming that the text prompt is uniform and interpreting as the cross-attention map intensity at the corresponding location, we have:
@@ 4.1.2 Few-shot
In the few-shot setup, normal reference images are utilized. For the -shot case, images are used to construct feature memory banks, which are then compared with the features of the query image.
@@
where is the set of blocks . That is, a score is computed based on the similarity, and the smallest score among all reference features in is conservatively used as the anomaly score for . Finally, the average over is taken as the final anomaly score.
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPRO | AUROC | AUPRO | ||
| 0-shot | Trans-MM [6] | 57.50.0 | 21.90.0 | 49.40.0 | 10.20.0 |
| MaskCLIP [57] | 63.70.0 | 40.50.0 | 60.90.0 | 27.30.0 | |
| WinCLIP [25] | 85.10.0 | 64.60.0 | 79.60.0 | 56.80.0 | |
| CLIPFUSION (ours) | 90.90.0 | 87.10.0 | 92.10.0 | 83.30.0 | |
| 1-shot | PaDiM [11] | 89.30.9 | 73.32.0 | 89.90.8 | 64.32.4 |
| PatchCore [43] | 92.01.0 | 79.72.0 | 95.40.6 | 80.52.5 | |
| WinCLIP [25] | 95.20.5 | 87.11.2 | 96.40.4 | 85.12.1 | |
| PromptAD [29] | 94.90.8 | 87.71.5 | 95.70.9 | 84.01.2 | |
| CLIPFUSION (ours) | 95.80.1 | 90.40.1 | 97.30.1 | 87.00.3 | |
| 2-shot | PaDiM [11] | 91.30.7 | 78.21.8 | 92.00.7 | 70.12.6 |
| PatchCore [43] | 93.30.6 | 82.31.3 | 96.10.5 | 82.62.3 | |
| WinCLIP [25] | 96.00.3 | 88.40.9 | 96.80.3 | 86.21.4 | |
| PromptAD [29] | 95.00.7 | 88.31.0 | 96.30.9 | 84.80.8 | |
| CLIPFUSION (ours) | 96.30.1 | 91.40.2 | 97.70.1 | 87.30.2 | |
| 4-shot | PaDiM [11] | 92.60.7 | 81.31.9 | 93.20.5 | 72.61.9 |
| PatchCore [43] | 94.30.5 | 84.31.6 | 96.80.3 | 84.91.4 | |
| WinCLIP [25] | 96.20.3 | 89.00.8 | 97.20.2 | 87.60.9 | |
| PromptAD [29] | 95.50.6 | 89.00.6 | 96.30.8 | 85.10.9 | |
| CLIPFUSION (ours) | 96.80.1 | 92.10.2 | 98.00.1 | 87.70.5 | |
@@
Reference images are passed through the diffusion denoiser to extract features, which are stored in the memory bank. The denoiser consists of three parts: an encoder (down-blocks), a bottleneck (mid-block), and a decoder (up-blocks). In the encoder, during compression, information is passed to the decoder via skip connections. The bottleneck is the most compressed stage, where much of the detailed information is lost, making it unsuitable for feature extraction. Therefore, we extract features from the decoder. Specifically, we focus on the middle and later decoder blocks, as they provide the most semantically rich features [1, 56].

### v3

@@ Abstract
Anomaly detection is a complex problem due to the ambiguity in defining anomalies, the diversity of anomaly types (e.g., local and global defect), and the scarcity of training data. As such, it necessitates a comprehensive model capable of capturing both low-level and high-level features, even with limited data. To address this, we propose CLIPFUSION, a method that leverages both discriminative and generative foundation models. Given the CLIP-based discriminative model’s limited capacity to capture fine-grained local details, we incorporate a diffusion-based generative model to complement its features. This integration yields a synergistic solution for anomaly detection. To this end, we propose using diffusion models as feature extractors for anomaly detection, and introduce carefully designed strategies to extract informative cross-attention and feature maps. Experimental results on benchmark datasets (MVTec-AD, VisA) demonstrate that CLIPFUSION consistently outperforms baseline methods in both anomaly segmentation and classification under both zero-shot and few-shot settings. We believe that our method underscores the effectiveness of multi-modal and multi-model fusion in tackling the multifaceted challenges of anomaly detection, providing a scalable solution for real-world applications.
@@ 1 Introduction
Anomaly detection is an important problem in real-world applications such as industry (bergmann2019mvtec; zou2022spot) and medicine (setio2017validation; menze2014multimodal). In this paper, we address anomaly classification and segmentation.111It is also referred to as anomaly detection and localization, respectively, in the literature. We use the terms anomaly classification and segmentation, and refer to anomaly detection as the overarching problem encompassing both. In anomaly classification (or segmentation), each image (or pixel) is classified as normal or anomalous.
@@
In practice, the number of abnormal images is small, while the potential types of abnormalities are highly diverse, making it challenging to obtain representative examples. Moreover, developing methods that can quickly adapt to new environments (e.g., new processes or products), even with a limited number of normal images, has become a critical research topic (rudolph2021same; sheynin2021hierarchical; zou2022spot; huang2022registration). To address this, we focus on the problem within the zero-shot or few-normal-shot222In this paper, ”few-shot” specifically refers to ”few-normal-shot”. setup.
@@
The recent advent of large-scale pretrained foundation models, e.g., GPT (brown2020language) and DALL-E (ramesh2021zero; ramesh2022hierarchical), has opened new horizons for addressing such challenges. The generalizability and scalability of foundation models make them particularly effective for tackling zero-shot and few-shot tasks such as anomaly detection. Effectively utilizing off-the-shelf foundation models is an active research direction. Our design follows this trend without retraining or introducing a shared module. Building on this, we present a novel methodology that combines the complementary strengths of discriminative and generative foundation models.
@@
Anomaly detection is a multifaceted problem, and various approaches have been developed based on different perspectives. First, anomaly detection has the aspect of one-class classification. This has led the community to methods that learn normality from normal images and judge what deviates from it as anomalous. Here, many methods extract normal features from normal images using a vision foundation model (cohen2020sub; defard2021padim; roth2022towards). Second, anomaly detection has the aspect of two-class classification. The community has employed vision-language foundation models, such as CLIP (radford2021learning), to compare a query image against normal and abnormal text prompts (jeong2023winclip; li2024promptad). Third, anomaly detection also has the aspect of semantic segmentation. Many anomalous features are fine-grained ones that appear in textures or edges. WinCLIP (window-based CLIP) attempted to solve this by sliding windows at multiple scales. However, it is known that CLIP has limitations in learning fine-grained features (wysoczanska2025clip; jiang2023clip), and such an approach might be suboptimal. Also, it requires hundreds of forward passes due to the exhaustive sliding of multi-scale windows.
@@
The multifaceted nature of anomaly detection makes it challenging to address using only discriminative models. So, we aim to create synergy by fusing a discriminative vision-language model (VLM) with a generative text-to-image (T2I) model. These are two fundamentally distinct approaches to aligning images and text. Discriminative models, such as CLIP, directly align embeddings in the same space, while generative models, like diffusion models, align them indirectly by jointly performing a common task. Diffusion models align images and text through a generative process that preserves fine-grained semantics. Consequently, they may capture more detailed visual patterns. Therefore, it can be beneficial to leverage both high-level information extracted from the CLIP-based model and low-level information extracted from the diffusion-based model for effective anomaly detection.
@@
We propose CLIPFUSION (= CLIP + Diffusion), a novel method that combines the strengths of multi-modal (language and vision) and multi-model (CLIP and diffusion) approaches. In the zero-shot scenario, CLIPFUSION effectively identifies anomalies by integrating image-text alignment from a CLIP-based model and cross-attention maps (chefer2023attend; zhang2023diffcloth; liu2024towards) from a diffusion-based model. In the few-shot scenario, it further enhances its performance by additionally leveraging feature maps from the CLIP image encoder and the diffusion denoiser, enabling a more comprehensive comparison between the query image and reference images. Figure 1 illustrates qualitative results, showcasing the model’s ability to capture fine-grained anomalies. CLIPFUSION exemplifies a holistic framework that synergizes complementary foundation models, and has the potential to extend beyond anomaly detection and advance other recognition tasks.
@@
We propose the use of diffusion models as feature extractors for anomaly detection. To this end, we exploit cross-attention and feature maps and carefully design extraction strategies. Since we extract intermediate features from an off-the-shelf diffusion model without any training or iterative denoising, our method is fast and efficient.
We introduce a novel approach to anomaly detection that jointly leverages features from CLIP and diffusion models. We demonstrate the complementary benefits of combining features from both models. Since diffusion models compensate for CLIP’s limitations in capturing fine-grained local details, our method eliminates the need for time-consuming techniques such as multi-scale sliding windows.
In terms of performance, our method consistently outperforms baselines. Unlike most prior works, it is also versatile in scope: it can handle both segmentation and classification tasks under zero-shot and few-shot settings.
Our method requires no additional training, which enables fast real-world deployment. Even in few-shot settings, the samples are used solely for feature extraction rather than training, making the method more robust. In contrast, training-based methods are sensitive to the choice of when to stop training, which is challenging in few-shot setups due to the absence of a validation set.
@@ 2 Related Work
In works that utilize vision(-language) models in anomaly detection, features are extracted from the intermediate layers of a model pretrained on a discriminative task (e.g., classification or contrastive learning). Anomaly detection is performed by comparing the features of the query image with the features of reference images. Some works utilize ResNet-based models pretrained on ImageNet in the few-shot regime. In SPADE (cohen2020sub), a -nearest neighbors algorithm is employed to identify the reference images closest to the query image, which are then used as the basis for anomaly detection. In PaDiM (defard2021padim), features extracted from reference images are modeled as Gaussian distributions, and the features of the query image are compared against these distributions. In PatchCore (roth2022towards), a memory bank is constructed using the features of reference images, and the features of a query image are directly compared against this memory bank. WinCLIP (jeong2023winclip) applies local windows and combines the results through harmonic aggregation. PromptAD (li2024promptad) learns prompt tokens through contrastive training.
A separate line of work assumes access to additional data and leverages it. Some methods require an auxiliary dataset containing both normal and anomaly images, e.g., training on the MVTec-AD dataset and inference on the VisA dataset (hu2024anomalydiffusion; zhu2024toward; yao2024resad; li2024one; cao2024adaclip; zhou2023anomalyclip; gu2024filo; qu2024vcp). AnomalyGPT (gu2024anomalygpt) relies on an auxiliary dataset with a large number of normal samples. In practice, acquiring such data with matching characteristics is challenging, and these methods often struggle when deployed on datasets with distribution shifts. Moreover, the need for additional training increases deployment time. Some methods also rely on ground-truth segmentation masks, which are difficult to obtain in real-world scenarios (hu2024anomalydiffusion; cao2024adaclip; zhou2023anomalyclip). MuSc (li2024musc) leverages the test set by comparing the features of test images, and is only feasible for batch-processing of the test set.
@@
Recently, there has been growing interest in applying diffusion models to discriminative tasks (fuest2024diffusion). In DDPM-Seg (baranchuk2021label), feature maps are constructed using activations extracted from intermediate layers of the denoiser and are utilized for semantic segmentation. VPD (zhao2023unleashing) further incorporates cross-attention maps for semantic segmentation and depth estimation. These methods require fine-tuning for downstream tasks. On the other hand, DiffusionClassifier (li2023your) performs zero-shot classification by determining which text conditioning best predicts the noise. P2PCAC (hertz2022prompt) and USCSD (hedlin2024unsupervised) reveal that cross-attention maps contain rich information about spatial layouts through image editing and semantic correspondence tasks.
@@ 3 Preliminaries
CLIP (radford2021learning) consists of an image encoder and a text encoder . These encoders map images and texts into the same embedding space. The model is trained to match correct image-text pairs with high similarity while assigning low similarity to incorrect pairs. This process is optimized using the following contrastive loss (chen2020simple):
@@
Diffusion
A diffusion model (ho2020denoising) is trained through a sequential process that involves forward diffusion and backward denoising. In the diffusion process, random noise is progressively added to a clean image over several steps, following the equation:
where is the noise schedule hyperparameter, which determines the amount of noise added to the image (nichol2021improved). In the denoising process, random noise is progressively removed over several steps using the following equation to restore the original image:
In this process, the denoiser learns to predict the added noise from the noisy image at a given time step (song2020score). During inference, random noise is sampled, and the denoiser iteratively removes it to generate a new image. By conditioning on a text prompt , the denoiser guides the generation process to align the resulting image with the provided text. Diffusion models demonstrate superior performance in understanding image details by generating high-quality images compared to other generative models (dhariwal2021diffusion). They can also be utilized for inpainting a specific region of an image. At each step, the untouched region is preserved by replacing it with the corresponding part of the original image, ensuring that only the specified region is generated and seamlessly blended with the rest of the image (rombach2021highresolution).
@@ 4 Method
The final model, CLIPFUSION, is a fusion of our CLIP-based and diffusion-based models. For convenience, we call the CLIP-based model PatchCLIP (patch-based CLIP) and the diffusion-based model MapDiff (map-based diffusion). We first outline our feature extraction strategies from both models, tailored for anomaly detection.
@@
CLIP processes an image by dividing it into multiple patches and outputs both patch embeddings and a class token embedding. While the class token embedding is aligned with text embeddings, patch embeddings can provide additional information, which we incorporate to enrich the spatial representation. Given that fine-grained details are handled by the diffusion-based model, PatchCLIP provides a simple yet effective spatial prior, thereby avoiding the cumbersome multi-scale window sliding in WinCLIP.
Additionally, we utilize features extracted from CLIP’s image encoder. To achieve this, we make use of the VV-attention mechanism proposed in CLIP surgery (li2023clip; li2024promptad). This mechanism was introduced to complement CLIP’s QK-attention mechanism, which is specialized in global feature extraction. It facilitates local feature extraction by enabling direct interactions between values (hence the name VV-attention) that carry information about the patches themselves.
@@
We extract and utilize cross-attention maps and feature maps from the diffusion denoiser. To achieve this, we use an inpainting pipeline that takes a clean image as input, rather than a generation pipeline that starts with random noise. In the inpainting pipeline, a mask is provided to specify the region where noise will be added. Here, we use the trick of applying an empty mask. This prevents noise from being added, but the denoiser still perceives the original image as noisy, thereby allowing us to control the timestep and extract features at the desired level of detail.
@@ 4.1.1 Zero-shot
In the zero-shot setup, given a query image, we leverage the relationships between images and texts learned by multi-modal models to create anomaly score maps.
@@
The query image is input to the CLIP image encoder (e.g., ViT), which divides it into patches, processes them, and outputs patch embeddings. On the other hand, normal and abnormal text prompts are passed to the CLIP text encoder, producing normal and abnormal text embeddings. Then, the degree (Equation 1) to which each patch embedding is aligned to the abnormal text embedding is calculated for all patches and organized into a map. This is upsampled to the original image size through interpolation to obtain .
@@
where [object] refers to the entity (e.g., “bottle”) in the image, and [state] describes a possible abnormal condition (e.g., “crack”) of it. The [state] token interacts with the image as it passes through the denoiser (e.g., U-Net), generating a cross-attention map. The cross-attention map of a diffusion model is a heatmap showing where and how strongly a text token is attended to across image locations. Because we use a token that denotes an anomalous state, the attention weight at a location serves as a proxy for anomaly likelihood. That is, a higher value at a specific location in the map indicates that the location responds more strongly to the [state] token (i.e., the abnormal condition). Therefore, we can use this cross-attention map as an anomaly score map (Figure 3). Additionally, we calculate the average of cross-attention maps derived using various possible abnormal conditions (e.g., “crack”, “hole”, “residue”, “damage”) to obtain .
@@
where represents the set of text prompts that describe abnormal conditions. Assuming that the text prompt is uniform and interpreting as the cross-attention map intensity at the corresponding location:
@@ 4.1.2 Few-shot
In the few-shot setup, normal reference images are utilized. For the -shot case, images are used to construct feature memory banks, which are then compared with the features of the query image. In general, we extract normal features from normal reference images by passing them through the CLIP image encoder and the diffusion denoiser. Then, the query image is passed through and the extracted features are compared with the normal features. The degree of deviation is used as an anomaly score.
@@
where is the set of blocks . That is, a score is computed based on the similarity, and the smallest among all reference features in is conservatively used as the anomaly score for . The final anomaly score is the average over .
@@
Reference images are passed through the diffusion denoiser to extract features, which are stored in the memory bank. The denoiser consists of three parts: an encoder (down-blocks), a bottleneck (mid-block), and a decoder (up-blocks). Information in the encoder is passed to the decoder via skip connections. The bottleneck is the most compressed stage, where much of the detailed information is lost, making it unsuitable for feature extraction. Therefore, we extract features from the decoder. Specifically, we focus on the middle and later decoder blocks, as they provide the most semantically rich features (baranchuk2021label; zhao2023unleashing).